\begin{thebibliography}{99}
		
%\bibitem{Aaron}
%J. Aaronson.
%\newblock An Introduction to Infinite Ergodic Theory (Mathematical Surveys and Monographs, 50).
%\newblock {\em American Mathematical Society}, Providence, RI, 1997.

\bibitem{A76}
G. Atkinson.
\newblock Recurrence of co-cycles and random walks.
\newblock {\em J. London Math. Soc.} 13(2): 486--488, 1976.


\bibitem{BJ21}
J.~Barral and X.~Jin
\newblock On the Action of Multiplicative Cascades on Measures.
\newblock {\em Int. Math. Res. Not.}, 2022(18): 13857--13896, 2022.


\bibitem{CF51}
K. L.~Chung and W. H. J.~Fuchs.
\newblock On the distribution of values of sums of random variables.
\newblock {\em Mem. Amer. Math. Soc.}, 6:1--12, 1951.

\bibitem{D82}
F. M.~Dekking.
\newblock On transience and recurrence of Generalized random walks.
\newblock {\em Z. Wahrsch. Verw. Gebiete}, 61:459--465, 1982.

\bibitem{Der}
Y.~Derriennic.
\newblock Ergodic theorem, reversibility and the filling scheme.
\newblock {\em Colloq. Math.}, 118(2):599--608, 2010.


%\bibitem{Dur}
%R. Durrett.
%\textit{Probability: Theory and Examples, 5th ed.} Cambridge Series in Statistical and Probabilistic Mathematics, 2019.

\bibitem{Ke}
H.~Kesten.
\newblock Sums of stationary sequences cannot grow slower than linearly.
\newblock {\em Proc. Amer. Math. Soc.}, 49:205--211, 1975.

%\bibitem{VO}
%M.~Viana and K.~Oliveira
%\newblock Foundations of Ergodic Theory (Cambridge Studies in Advanced Mathematics).
%\newblock Cambridge: Cambridge University Press, 2016.

\end{thebibliography}

\documentclass[12pt]{amsart}

\usepackage[a4paper,left=2.5cm,right=2.5cm,top=2.5cm,bottom=2.5cm]{geometry}
\usepackage{amsmath}

\usepackage{graphicx}
\usepackage{amssymb}
\usepackage{bm}
\usepackage{enumerate}
\usepackage{setspace}
\usepackage{color}
\usepackage[colorinlistoftodos,textwidth=2cm]{todonotes}
\usepackage{tikz}
\usetikzlibrary{patterns,patterns.meta}

\newtheorem{theorem}{Theorem}[section]
\newtheorem{corollary}[theorem]{Corollary}

\newtheorem{remark}[theorem]{Remark}
\numberwithin{equation}{section}

\begin{document}

\title{A Chung-Fuchs type theorem for skew product dynamical systems}

\author{Xiong Jin}
\address{Department of Mathematics, University of Manchester, Oxford Road, Manchester M13 9PL, United Kingdom}
\email{xiong.jin@manchester.ac.uk}

\begin{abstract}
We prove a Chung-Fuchs type theorem for skew product dynamical systems such that for a measurable function on such a system, if its Birkhoff average converges to zero almost surely, and on typical fibres its Birkhoff sums have a non-trivial independent structure, then its associated generalised random walk oscillates, that is the supremum of the random walk equals to $+\infty$ and the infimum equals to $-\infty$.
\end{abstract}

\maketitle

\section{Introduction}

Let $(\Omega,\mathcal{A},\mathbb{P})$ be a probability space. Let $\{W_n\}_{n\ge 1}$ be a sequence of independent and identically distributed (i.i.d.) random variables in $\mathbb{R}$. Let 
\[
S=\{S_n=W_1+\cdots +W_n\}_{n\ge 1}
\]
denote its associated random walk. We say that $S$ is recurrent if for all $c>0$,
\[
\mathbb{P}(|S_n|\le c \text{ for infinitely many } n)=1.
\]
Regarding the recurrence of random walks we have the following well-known theorem of Chung and Fuchs \cite{CF51}:
\begin{theorem}[Chung and Fuchs 1951]
If $\mathbb{E}(|W_1|)<\infty$ and $\mathbb{E}(W_1)=0$, then $S$ is recurrent. Furthermore, if $\mathbb{P}(W_1\neq 0)>0$, then $S$ oscillates, that is
\begin{equation}\label{osc}
\mathbb{P}\big(\liminf_{n\to\infty} S_n=-\infty\big)=\mathbb{P}\big(\limsup_{n\to\infty} S_n=+\infty\big)=1.
\end{equation}
\end{theorem}
In \cite{CF51}, the condition ``$\mathbb{E}(|W_1|)<\infty$ and $\mathbb{E}(W_1)=0$" is only to ensure that, by the strong law of large numbers,
\begin{equation}\label{sll}
\mathbb{P}\Big(\lim_{n\to\infty} \frac{S_n}{n}=0\Big)=1.
\end{equation}
So Chung and Fuchs' theorem can be stated in a slightly more general way without $L^1$-integrable assumption that
\[
\eqref{sll} \ \Rightarrow \ S \text{ is recurrent}.
\]
For  \eqref{osc}, Chung and Fuchs showed that the set of recurrent points is a closed additive subgroup of $\mathbb{R}$, hence it is either empty, or the whole space $\mathbb{R}$, or a lattice $a\cdot\mathbb{Z}$ for some $a\in\mathbb{R}$. So if $S$ is recurrent, then the set of recurrent points is non-empty, thus it is either $\mathbb{R}$ or $a\cdot\mathbb{Z}$ for some $a\neq 0$ (since $\mathbb{P}(W_1\neq 0)>0$), and either case implies \eqref{osc}.

One may consider the Chung-Fuchs theorem for more general random walks: Let $(X,\mathcal{B},T,\mu)$ be a measure-preserving dynamical system (m.p.d.s) with $\mu(X)=1$. Let $f:X\to\mathbb{R}$ be a measurable function. Denote by
\[
S_f=\{S_nf=f+f\circ T+\cdots +f\circ T^{n-1}\}_{n\ge 1}
\]
its associated sequence of Birkhoff sums, or in other words, its associated generalised random walk. We say that $S_f$ is recurrent if for all $c>0$,
\[
\mu(|S_nf|\le c \text{ for infinitely many } n)=1.
\]
The following extension of Chung-Fuchs theorem to generalised random walks is essentially due to Dekking \cite{D82} (see also \cite{A76} when assuming $f\in L^1(\mu)$).

\begin{theorem}[Dekking 1982]\label{Dekking}
If $\mu$ is ergodic and
\begin{equation}\label{bet}
\mu\Big(\lim_{n\to\infty} \frac{S_nf}{n}=0\Big)=1,
\end{equation}
then $S_f$ is recurrent.
\end{theorem}

For generalised random walks, the statement that \eqref{bet} plus $f\not\equiv 0$ implies \eqref{osc} is no longer true. A simple counter example is the following: Let $X=\{x_0,x_1\}$ be a space of two points and let the transformation $T$ be $T(x_0)=x_1$, $T(x_1)=x_0$ and let $\mu(\{x_0\})=\mu(\{x_1\})=\frac{1}{2}$. Take $f:X\to\mathbb{R}$ to be $f(x_0)=1$, $f(x_1)=-1$. Then $S_nf$ is periodic and \eqref{bet} holds, but both liminf and limsup of $S_nf$ are bounded.

\medskip

It is natural to ask for which settings between random walks and generalised random walks it holds that \eqref{bet} plus $f\not\equiv 0$ implies \eqref{osc}. In \cite{BJ21} the authors studied the action of Mandelbrot cascades on ergodic measures, and in the estimation of the degeneracy of this action it was sufficient to show that
\begin{equation}\label{ui}
\mathbb{Q}\Big(\limsup_{n\to\infty} S_nf+S_nF=+\infty\Big)=1,
\end{equation}
where, in brief, the probability measure $\mathbb{Q}$ is a skew-product extension of the ergodic measure $\mu$ to $X\times\Omega$, $S_nf(\cdot)$ is a generalised random walk under $\mu$ and $S_nF(x,\cdot)$ is a random walk under $\mathbb{Q}_x$ for $\mu$-a.e. $x$, where the probability measure $\mathbb{Q}_x$ on $\Omega$ is the disintegration of $\mathbb{Q}$ w.r.t. $\mu$. In the sub-critical case when $S_nf+S_nF$ has a positive drift, that is $\mathbb{Q}$-a.s. $\lim_{n\to\infty}\frac{S_nf+S_nF}{n}=c>0$, it is easy to deduce \eqref{ui}. The critical case when $\mathbb{Q}(\lim_{n\to\infty}\frac{S_nf+S_nF}{n}=0)=1$ corresponds to \eqref{bet} and it becomes more delicate to verify \eqref{ui}, which has led us to derive the following result (a  more general and more abstract version of the result in \cite{BJ21}).

\subsection*{Main result}

Recall that $(X,\mathcal{B},T,\mu)$ is a m.p.d.s with $\mu(X)=1$. Consider a skew-product measure-preserving dynamical system $(X\times\Omega,\hat{\mathcal{B}},\hat{T},\mathbb{Q})$ where
\begin{itemize}
\item[-] $\hat{\mathcal{B}}=\mathcal{B}\otimes \mathcal{A}$.
\item[-] $\hat{T}(x,\omega)=(T(x),g_x(\omega))$ is a skew-product transformation, where for $x\in X$, $g_x:\Omega\to\Omega$ is $\mathcal{A}$-measurable and for $A\in\mathcal{A}$, $x\to g_x(A)$ is $\mathcal{B}$-measurable.
\item[-] $\Pi_X^*(\mathbb{Q})=\mu$, where $\Pi_X$ is the projection from $X\times\Omega$ onto $X$.
\end{itemize}
Consider the measurable partition
\[
\eta=\{\Pi_X^{-1}(x)=\{x\}\times\Omega:x\in X\}
\]
and let $\mathbb{Q}^\eta_{x,\omega}$ denote the conditional measure of $\mathbb{Q}$ w.r.t. $\eta$. For $\mathbb{Q}$-a.e. $(x,\omega)$, $\mathbb{Q}^\eta_{x,\omega}$ is a probability measure carried by $\{x\}\times\Omega$. Note that $\mathbb{Q}^\eta_{x,\omega}$ only depends on $x$. Hence, denoting $\Pi_\Omega$ the projection from $X\times\Omega$ onto $\Omega$, we may define a family of probability measures $\mathbb{Q}_x=\Pi_\Omega^*(\mathbb{Q}^\eta_{x,\omega})$ on $\Omega$ for $\mu$-a.e. $x\in X$ such that for any measurable function $F:X\times\Omega\to\mathbb{R}$,
\[
\int_{X\times\Omega} F(x,\omega)\,\mathbb{Q}(\mathrm{d} (x,\omega))=\int_X\int_\Omega F(x,\omega)\,\mathbb{Q}_x(\mathrm{d}\omega)\mu(\mathrm{d}x).
\]
The family of probability measures $\mathbb{Q}_x$ for $\mu$-a.e. $x\in X$ is also referred as the disintegration of $\mathbb{Q}$ with respect to $\mu$.

Let $F:X\times\Omega\to\mathbb{R}$ be a measurable function such that
\begin{itemize}
\item[(A)] for $\mu$-a.e. $x\in X$,
\subitem (i) $F(x,\cdot)$ is not $\mathbb{Q}_x$-a.s. a constant;
\subitem (ii) the sequence $\{F(x,\cdot),F\circ \hat{T}(x,\cdot),F\circ \hat{T}^2(x,\cdot),\cdots\}$ is an independent sequence of random variables under $\mathbb{Q}_x$.
\end{itemize}
For $n\ge 1$ denote by
\[
\hat{S}_nF=F+F\circ \hat{T}+\cdots+F\circ \hat{T}^{n-1}.
\]
We have the following extension of Chung and Fuchs theorem.
\begin{theorem}\label{mainthm}
Assume (A).  If
\[
\mathbb{Q}\Big(\lim_{n\to\infty} \frac{\hat{S}_nF}{n}=0\Big)=1,
\]
then
\[
\mathbb{Q}\Big(\liminf_{n\to\infty} \hat{S}_nF=-\infty\Big)>0  \text{ and }  \mathbb{Q}\Big(\limsup_{n\to\infty} \hat{S}_nF=+\infty\Big)>0.
\]
\end{theorem}

\begin{remark}
If we further assume that $\mathbb{Q}$ is $\hat{T}$-ergodic, then, since $\{\liminf_{n\to\infty} \hat{S}_nF=-\infty\}$ and $\{\limsup_{n\to\infty} \hat{S}_nF=+\infty\}$ are $\hat{T}$-invariant sets, Theorem \ref{mainthm} implies
\[
\mathbb{Q}\Big( \liminf_{n\to\infty} \hat{S}_nF=-\infty\Big)=\mathbb{Q}\Big( \limsup_{n\to\infty} \hat{S}_nF=+\infty\Big)=1.
\]
\end{remark}

\begin{remark}
Assumption (A) holds for a more general class of skew-product measures $\mathbb{Q}$ than the Peyri\`ere measure studied in \cite{BJ21}. In particular we are not requiring that the law of $F(x,\cdot)$ under $\mathbb{Q}_x$ to be independent of $x$.
\end{remark}

\begin{remark}
We improved the proofs comparing to \cite{BJ21}. Firstly we used a result of Kesten to remove the method of using the stopping time that $F\circ\hat{T}^n(x,\cdot)$ firstly becomes positive, then we added an induced system argument to deduce that a measurable (not necessarily integrable) coboundary $u-u\circ T=v$ with $u,v\ge 0$ is equal to zero almost everywhere. 
\end{remark}

\subsection*{Example}

We may consider the following canonical example: let $(X\times Y, \mathcal{B}_{X\times Y},\tilde{T}, \Theta)$ be a skew product dynamical system of $(X,\mathcal{B},T,\mu)$, where the skew product $\tilde{T}(x,y)=(T(x),\varphi_x(y))$ is defined for a family of measurable functions $\varphi_x:Y\to Y$ and the $\tilde{T}$-invariant measure $\Theta$ satisfies $\Pi_X^*(\Theta)=\mu$. For $\mu$-a.e. $x\in X$ let $\Theta_x$ denote the disintegration of $\Theta$ w.r.t. $\mu$. We define the skew-product dynamical system $(X\times\Omega,\hat{\mathcal{B}},\hat{T},\mathbb{Q})$ as follows: Let $\Omega$ be the infinite product space of $Y$, that is, 
\[
\Omega=\bigotimes_{\mathbb{N}} Y.
\]
The family of measurable functions $g_x:\Omega\to\Omega$ used to define the skew product $\hat{T}(x,\omega)=(T(x),g_x(\omega))$ is taken to be the left-shift operator $\sigma$, that is,
\[
g_x(\omega)=\sigma(\omega)=(y_2,y_3,\cdots) \text{ for } \omega=(y_n)_{n\ge 1}\in \Omega.
\]
The $\hat{T}$-invariant measure $\mathbb{Q}$ on $X\times\Omega$ is defined as
\[
\mathbb{Q}(\mathrm{d}x,\mathrm{d}\omega)=\mu(\mathrm{d}x)\mathbb{Q}_x(\mathrm{d}\omega),
\]
where the fibre measures $\mathbb{Q}_x$ for $\mu$-a.e. $x\in X$ is the product measure
\[
\mathbb{Q}_x(\mathrm{d}\omega)=\prod_{k=1}^\infty\Theta_{T^{k-1}(x)}(\mathrm{d}y_k) \text{ for } \omega=(y_n)_{n\in\mathbb{N}}\in \Omega.
\]
Let $h:X\times Y\mapsto \mathbb{R}$ be a measurable function. Define the measurable function $F:X\times \Omega\to\mathbb{R}$ by
\[
F(x,\omega)=h(x,y_1) \text{ for } x\in X \text{ and } \omega=(y_n)_{n\in\mathbb{N}}\in\Omega.
\]
Then for $k\ge 1$ we have
\[
F\circ \hat{T}^{k-1}(x,\omega)=h(T^{k-1}(x),y_k) \text{ for } x\in X \text{ and } \omega=(y_n)_{n\in\mathbb{N}}\in\Omega.
\]
By the product structure it is straightforward that the sequence of random variables 
\[
\{F(x,\cdot),F\circ \hat{T}(x,\cdot),F\circ \hat{T}^2(x,\cdot),\cdots\}
\]
is independent under $\mathbb{Q}_x$, therefore in this setting we have the following corollary from Theorem \ref{mainthm} (for simplicity we assume $h$ to be in $L^1$ and $\mathbb{Q}$ is ergodic).

\begin{corollary}
Assume that
\begin{itemize}
\item for $\mu$-a.e. $x\in X$, $h(x,\cdot)$ is not $\Theta_x$-a.s. a constant;
\item $h\in L^1(\Theta)$, $\int_{X\times Y} h\,\mathrm{d}\Theta=0$ and $\mathbb{Q}$ is $\hat{T}$-ergodic.
\end{itemize}
Then for $\mu$-a.e. $x\in X$ and for $\prod_{k=1}^\infty\Theta_{T^{k-1}(x)}$-a.e. $(y_k)_{k\in \mathbb{N}}\in \bigotimes_{\mathbb{N}} Y$,
\[
\liminf_{n\to\infty} \sum_{k=1}^n h(T^{k-1}(x),y_k)=-\infty \text{ and }
\limsup_{n\to\infty} \sum_{k=1}^n h(T^{k-1}(x),y_k)=+\infty.
\]
\end{corollary}

\section{Proof of Theorem \ref{mainthm}}\label{proofs}

We shall use the filling scheme, see \cite{Der} for example.

For $a\in \mathbb{R}$ denote by $a^+=\max\{a,0\}$ and $a^-=\max\{-a,0\}$. 

For $n\ge 1$ define
\[
G_n:=\max_{1\le k \le n} \hat{S}_kF.
\]
We have
\[
G_{n+1}^+-G_{n+1}^-=G_{n+1}=F+G_n^+\circ \hat{T}.
\]
Let $G=\lim_{n\to\infty} G_n$. Then
\begin{equation}\label{Gpm}
G^+-G^-=F+G^+\circ \hat{T}.
\end{equation}
From the fact 
\[
\mathbb{Q}\Big(\lim_{n\to\infty} \frac{\hat{S}_nF}{n}=0\Big)=1
\]
we deduce that, $\mathbb{Q}$-a.s., $\hat{S}_nF$ changes signs in the wide sense infinite many times, that is, $\hat{S}_nF>0$ or $\hat{S}_nF<0$ cannot hold for all $n$ large enough. This is due to a result of Kesten \cite{Ke} that the sums of stationary sequences cannot grow slower than linearly, see \cite[Section 5.(c)]{Der} for example. This implies that, $\mathbb{Q}$-a.s., $G\ge 0$ hence $G^-=0$, thus we may write from \eqref{Gpm} that, $\mathbb{Q}$-a.s.,
\begin{equation}\label{coB}
G^+=F+G^+\circ \hat{T}.
\end{equation}

Assume that $\mathbb{Q}(G^+<\infty)=1$. This implies that for $\mu$-a.e. $x\in X$,
\[
\mathbb{Q}_x(G^+(x,\cdot)<\infty)=1.
\]

Fix $t\in\mathbb{R}\setminus \{0\}$. For $\mu$-a.e. $x\in X$ denote by
\[
\Phi_t(x):=\mathbb{E}_{\mathbb{Q}_{x}}(e^{it G^+(x,\cdot)}) \text{ and } \phi_t(x):=\mathbb{E}_{\mathbb{Q}_{x}}(e^{it F(x,\cdot)}).
\]
Then by \eqref{coB} and (A)(ii) we have, for $\mu$-a.e. $x\in X$,
\[
\Phi_t(x)=\phi_t(x)\times \Phi_t\circ T(x).
\]
Write $u(x)=-\log |\Phi_t(x)|$ and $v(x)=-\log |\phi_t(x)|$ with the convention that $-\log 0=\infty$. Then we have, $\mu$-a.s., $u,v\ge 0$ and
\[
u=v+u\circ T.
\]
Suppose that $\mu(u<\infty)>0$. Take $N>0$ large enough such that
\[
X_N=X\cap \{u\le N\}
\]
has positive $\mu$-mass. Let $(X_N,\mathcal{B}_N,T_N,\mu_N)$ denote the induced dynamical system of $(X,\mathcal{B},T,\mu)$ by $X_N$, where $\mathcal{B}_N$ is $\mathcal{B}$ restricted to $X_N$, $\mu_N=\frac{1}{\mu(X_N)}\mu|_{X_N}$ is the normalised $\mu$ restricted to $X_N$, and, denoting $\rho_N(x):=\inf\{n\ge 1:T^n(x)\in X_N\}$ the first visiting time to $X_N$, $T_N(x)=T^{\rho_{X_N}(x)}(x)$ is the induced transform from $X_N$ to $X_N$.

On $(X_N,\mathcal{B}_N,T_N,\mu_N)$ we obtain the co-boundary equation
\[
u-u\circ T_N=v_N,
\]
where
\[
v_N(x)=v(x)+v\circ T(x)+\cdots +v\circ T^{\rho_{X_N}(x)-1}(x).
\]
On $X_N$ we have $0\le u\le N$, $0\le u\circ T_N\le N$ and $0\le v_N\le u\le N$ are all bounded functions hence integrable. Thus by the $T_N$-invariance of $\mu_N$ and Birkhorff ergodic theorem we deduce that
\[
\mathbb{E}_{\mu_N}( v_N \,|\, \mathcal{I}_N)(x)=0
\]
for $\mu_N$-a.e. $x\in X_N$, where $\mathcal{I}_N$ is the $T_N$-invariant $\sigma$-algebra of $X_N$. Since $v$ is non-negative, $v_N$ is also nonnegative, we deduce that $v_N(x)=0$ and therefore $v(x)=0$ for $\mu_N$-a.e. $x\in X_N$. Since $N$ is arbitrary, we deduce that $v(x)=0$ or equivalently $|\phi_t(x)|=1$ for $\mu$-a.e. $x\in X\cap \{u<\infty\}$. Note that $|\phi_t(x)|=1$ implies $F(x,\cdot)$ is $\mathbb{Q}_x$-a.s. a constant, which is contradict to (A)(i). Therefore $\mu(u<\infty)=0$, or in other words,
\begin{equation}\label{m0}
\mu(x\in X: \Phi_t(x)=0)=1.
\end{equation}

We may take a countable sequence $t_n$ tending to $0$. Then \eqref{m0} implies that
\[
\mu(x\in X: \Phi_{t_n}(x)=0 \text{ for all } n \ge 1)=1.
\]
But this is not possible since for $\mu$-a.e. $x\in X$, $\mathbb{Q}_x(G^+(x,\cdot)<\infty)=1$, that is, $G^+(x,\cdot)$ is a proper random variable under $\mathbb{Q}_x$, thus $\mathbb{E}_{\mathbb{Q}_{x}}(e^{it G^+(x,\cdot)})$ is continuous in $t$, hence non-vanishing around $0$.

We finally deduce that $\mathbb{Q}(G^+<\infty)<1$, or in other words,
\[
\mathbb{Q}\Big(\limsup_{n\to\infty} \hat{S}_nF=+\infty\Big)>0.
\]
Since same arguments apply to $-F$, we also get
\[
\mathbb{Q}\Big(\liminf_{n\to\infty} \hat{S}_nF=-\infty\Big)>0.
\]

\end{document}